OpenStudy (anonymous):

Use a system of equations to solve the word problem. A rectangular table has a perimeter of 18 feet. Its length is 5 feet greater than its width. Find the dimensions. A. l = 8 and w = 3 B. l = 7 and w = 2 C. l = 2 and w = 7 D. l = 3 and w = 8

OpenStudy (snowfire):

Systems of equations relate two equations and allow us to see multiple separate relationships. In this case, can you tell me what the two equations are?

OpenStudy (anonymous):

18 and 5

OpenStudy (snowfire):

Those aren't equations. Start with the "18". That is the perimeter, right? And how do we define the perimeter of something?

OpenStudy (anonymous):

idk how to do this

OpenStudy (snowfire):

Giving up won't solve the problem. You need to take it piece by piece, so it doesn't overwhelm you. If I had a rectangle, and we used x and y to label its two sets of sides (because rectangles have two sets of sides that are equal in length), then what is its perimeter?

OpenStudy (anonymous):

do i add or what?

OpenStudy (snowfire):



OpenStudy (snowfire):

What is the perimeter of this rectangle?

OpenStudy (anonymous):

18+18=36 5+5=10 taotal =46

OpenStudy (snowfire):

The perimeter is 18. Use the variables I created in the picture as the numbers and set them equal to 18.

OpenStudy (anonymous):

I don't understand

OpenStudy (snowfire):

The perimeter of this rectangle is x+x+y+y=18, which is just 2x+2y=18

OpenStudy (anonymous):

yes

OpenStudy (snowfire):

Now we arbitrarily take either x or y to be the "length", for the purpose of utilizing the second part of the problem. Let's say x is the length. What is the relationship between length and width? (it is given in the problem)

OpenStudy (anonymous):

length is 18 and the width is 5

OpenStudy (snowfire):

Read the problem more carefully, that's not what it says.

OpenStudy (anonymous):

oh i mean lenght is 5 and width is 18

OpenStudy (snowfire):

Read again. 18 is the perimeter.

OpenStudy (anonymous):

oh....